OpenStudy (anonymous):

what is the 10th term of the sequence 64,16,4 a. 1/1024 b. 1/256 c. 1/4096 d. 1/496

OpenStudy (anonymous):

pretty sure its c

OpenStudy (anonymous):

keep dividing by four

OpenStudy (anonymous):

thanks so its c for sure?

OpenStudy (anonymous):

64, 16, 4, 1, 1/4, 1/16, 1/64, 1/256, 1/1024, 1/4096

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

notice, how did they get from 64 to 16, and how did they then get from 16 to 4, well\[\frac{ 64 }{ 4 }=16\]\[\frac{ 16 }{ 4 }=4\]so that means every number is being divided by 4, you could also notice that\[\frac{64}{4^0} \rightarrow\frac{ 64 }{ 4^1 } \rightarrow \frac{ 64 }{ 4^2 } \rightarrow \frac{ 64 }{ 4^3 }\] is the sequence that you have,

OpenStudy (anonymous):

by following this pattern, you'll get that the 10th number in the sequence is\[\frac{ 64 }{ 4^9 }=\frac{ 1 }{ 4096 }\]